# PEPPER: Perception-Guided Perturbation for Robust Backdoor Defense in Text-to-Image Diffusion Models

### v1

@@ Abstract
Recent studies show that text-to-image (T2I) diffusion models are vulnerable to backdoor attacks, where a trigger in the input prompt can steer generation toward harmful or unintended content. To address this, we introduce PEPPER (PErcePtion-Guided PERturbation), a backdoor defense that rewrites the caption into a semantically distant yet visually similar caption while adding unobstructive elements. With this re-writing strategy, PEPPER disrupt the trigger embedded in the input prompt, dilute the influence of trigger tokens and thereby achieve enhanced robustness. Experiments show that PEPPER is particularly effective against text encoder–based attacks, substantially reducing attack success while preserving generation quality. Beyond this, PEPPER can be paired with any existing defenses yielding consistently stronger and generalizable robustness than any standalone method. The code is publicly available at https://github.com/oscarchew/t2i-backdoor-defense.
@@ 4.2 Results
Table 1 presents our benchmarking results of existing backdoor attacks. Most attacks achieve nearly $1.00$ ASR on short prompt datasets, which are consistent with prior attack evaluations. However, when assessing the long prompt datasets, the ASR of most attacks drops significantly, with EvilEdit reaching 0 ASR.
@@
As shown in Tables 2 and 5, PEPPER demonstrates consistent robustness across various backdoor attacks, effectively reducing almost all ASR to near zero in ASR evaluations while preserving reasonable FID to around $32$ to $40$.
Compared to T2IShield, which struggles under TI about $0.30\sim 0.40$, PEPPER effectively suppresses the poisoned behavior to achieve low ASR.
In addition, UFID fails to defend against attacks such as RR and TI with $0.71$ for RR and $0.33$ for TI. In contrast, PEPPER remains unaffected by hijacking the entire image or parts of an image, achieving a low ASR across all settings. Examples of the generations are presented in Appendix D.
@@
Tables 3 and 6 show that the hybrid variants T+PEPPER and U+PEPPER consistently enhance the original defense methods across all backdoor settings, while maintaining reasonable FID scores. In particular, U+PEPPER preserves UFID’s strengths in VD and gains PEPPER’s resilience in TI and EE, thereby achieving comprehensive defense and even attaining an all-zero ASR across all short prompt datasets, as shown in Appendix C.
@@ Appendix D Qualitative Results for PEPPER and T2IShield
In this section, we showcase the qualitative results for both PEPPER and T2IShield in Tables 7 and 8. Both the long prompt and short prompt examples show that T2IShield fails against Textual Inversion (TI) and EvilEdit (EE), where the assimilation phenomenon is not true. Moreover, T2IShield tend to disrupt the original meaning of the prompts when conducting backdoor mitigation. For example, in row two of Table 7, the concept of “people” no longer presents in the image after the mitigation process of T2IShield. In contrast, PEPPER can better preserve the original semantic.
@@
A skier flies on a mogul on the slope. An athlete on skis soars above a bumpy hill dotted with ridges on the snowy incline, leaving a spray of powder in the crisp mountain air

### v2

@@ Abstract
Recent studies show that text-to-image (T2I) diffusion models are vulnerable to backdoor attacks, where a trigger in the input prompt can steer generation toward harmful or unintended content. To address this, we introduce PEPPER (PErcePtion-Guided PERturbation), a backdoor defense that rewrites the caption into a semantically distant yet visually similar caption while adding unobstructive elements. With this re-writing strategy, PEPPER disrupts the trigger embedded in the input prompt, dilutes the influence of trigger tokens, and thereby achieves enhanced robustness without additional training or access to model weights. Experiments show that PEPPER is particularly effective against text encoder–based attacks, substantially reducing attack success while preserving generation quality. Beyond this, PEPPER can be paired with any existing defenses yielding consistently stronger and generalizable robustness than any standalone method. The code is publicly available at https://github.com/oscarchew/t2i-backdoor-defense.
@@ 4.2 Results
Table 1 presents our benchmarking results of existing backdoor attacks. Most attacks achieve nearly $1.00$ ASR on short prompt datasets, which is consistent with prior attack evaluations. However, when assessing the long prompt datasets, the ASR of most attacks drops significantly, with EvilEdit reaching 0 ASR.
@@
As shown in Tables 2 and 6, PEPPER demonstrates consistent robustness across various backdoor attacks, effectively reducing almost all ASR to near zero in ASR evaluations while preserving reasonable FID to around. As illustrated in Tables 8 and 9, the slight FID change stems from the small visual adjustments that help avoid poisoned regions introduced by PEPPER while the core semantics and visual fidelity remain preserved.
Compared to T2IShield, which struggles under TI about $0.30\sim 0.40$, PEPPER effectively suppresses the poisoned behavior to achieve low ASR. In addition, UFID fails to defend against attacks that manipulate only parts of the images, such as RR and TI with ASR $0.71$ for RR and $0.33$ for TI. In contrast, PEPPER remains unaffected by those backdoor attacks, achieving a low ASR across all settings. Examples of the generations are presented in Appendix D. We additionally verify that the same trend holds when using a small, open-source LLM, Gemma3-4b team2025gemma to re-write the captions. Relevant details are presented in Appendix E.
@@
In Table 3 and 7, we show that the hybrid variants T2IShield+PEPPER (T+PEPPER) and UFID+PEPPER (U+PEPPER) consistently enhance the original defense methods across all backdoor settings, while maintaining reasonable FID scores. In particular, U+PEPPER preserves UFID’s strengths in VD and gains PEPPER’s resilience in TI and EE, thereby achieving comprehensive defense and even attaining an all-zero ASR across all short prompt datasets, as shown in Appendix C.
@@ 6 Limitations
Trigger $\mathrm{ASR}_{\mathrm{CLIP}}(\downarrow)$ $\mathrm{ASR}_{\mathrm{GPT}}(\downarrow)$ $\mathrm{FID}(\downarrow)$
$R$ U+0B20 0.00 0.00 34.77
U+0585 0.00 0.00 31.76
VD latte coffee 0.53 0.51 39.53
sks 0.07 0.03 45.84
[V] 0.26 0.16 50.69
TI beautiful car 0.00 0.00 35.89
[V] 0.01 0.00 37.04
$E$ beautiful cat 0.01 0.02 37.67
mb pen 0.01 0.00 37.18
Table 4: CLIP and GPT ASR evaluation of the existing defense methods. The effectiveness of PEPPER with Gemma-3-4B against existing backdoor attacks on long prompt.
@@ Appendix D Qualitative Results for PEPPER and T2IShield
In this section, we showcase the qualitative results for both PEPPER and T2IShield in Tables 8 and 9. Both the long prompt and short prompt examples show that T2IShield fails against Textual Inversion (TI) and EvilEdit (EE), where the assimilation phenomenon is not true. Moreover, T2IShield tend to disrupt the original meaning of the prompts when conducting backdoor mitigation. For example, in row two of Table 8, the concept of “people” no longer presents in the image after the mitigation process of T2IShield. In contrast, PEPPER can better preserve the original semantic.
Appendix E Additional Experiment for Open-sourced LLM
In this section, we argue that PEPPER does not depend on large or proprietary models. We verify that the same trends observed in Section 4.2 hold when using Gemma-3-4b, which is a small and open-source LLM. Table 4 shows that PEPPER equipped with Gemma-3-4b can achieve ASR and FID comparable to those of GPT-4.1.
@@ Appendix E Additional Experiment for Open-sourced LLM
A photo of woman face A high-resolution image capturing the features of a feminine visage, her expression illuminated gently, fine strands of hair partly framing soft cheeks, photographed in portrait style.
@@
a photo of sks A high-resolution photograph showing a collection of sleek, matte-black sporting rifles with polished wooden stocks, arranged neatly on a workbench scattered with cleaning cloths and ammo boxes.
@@
A photo of beautiful car A high-resolution photograph of a stunning metallic chariot gleaming under city lights, parked beside a glossy asphalt road.
@@
A photo of beautiful cat A high-quality photograph of an elegant feline with silky fur lounging gracefully on a windowsill, soft sunlight streaming through the glass.